OpenStudy (anonymous):

An outside circular ring has a circumference of 120 cm. What is the circumference of an inner ring which is 10 cm shorter in radius? Both circles have the same center. 35.2 cm 42.2 cm 57.2 cm 63.2 cm

OpenStudy (anonymous):

@RD⁴²

OpenStudy (anonymous):



OpenStudy (anonymous):

??

OpenStudy (anonymous):

Circumference of the outside circular ring: C = 2\(\pi\)r = 120 According to the question, inner circular ring is 10 cm shorter. so C = 2\(\pi\)(r-10). First, find r from outside circular ring. 2\(\pi\)r = 120 \(\pi\)r = 60 \(\boxed{r = \dfrac{60}{\pi}}\) Now plug it into inner circular ring equation. C = 2\(\pi\)\( \left( \dfrac{60}{\pi} - 10\right)\) Can you solve it now?

OpenStudy (anonymous):

so 60/3.14 - 10?

OpenStudy (anonymous):

All of this times 2\(\pi\)

OpenStudy (anonymous):

C = 2π( 60/π − 10 )

OpenStudy (anonymous):

so i got 9.1082802548 but then i need to times 2 * 3.14?

OpenStudy (anonymous):

i got 57.2

OpenStudy (anonymous):

cm

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

I tagged you in another question i need help with